If the moment of inertia of a tire is 0. 63 kg.m² and its weight is 24.5 N, what is the diameter of the tire? If the weight is doubled, what will be the moment of inertia?


1
Expert's answer
2021-12-03T12:44:21-0500

"I=mr^2"

"r=\sqrt{I/m}=\sqrt{0.63/(24.5/g)}=0.50" m

diameter of the tire:

"d=2r=2\cdot0.50=1" m


If the weight is doubled:

"I=2mr^2=2\cdot24.5\cdot0.5^2/g=1.25" kg"\cdot"